# Simple and Effective Specialized Representations for Fair Classifiers

### v1

@@ Appendix A Appendix
  • Section A.1: Complete derivations for Theorems 5.1, 5.2, and 5.3.
  • Section A.2: Analysis of MMD, its relation to characteristic-function distances, and links to adversarial representation learning.
  • Section A.3: Discussion of full moment-matching, the Russian-Roulette estimator for unbiased series truncation, and practical considerations.
  • Section A.4: Formal definitions and commentary on Demographic Parity, Equalized Odds, and other group-fairness criteria.
  • Section A.5: Extended comparisons against baselines, Pareto-front ablations over latent dimensions, and robustness checks.
  • Section A.6: Detailed descriptions of all benchmark datasets, preprocessing pipelines, and protected attribute encoding.
  • Section A.7: Implementation specifics, optimizer settings, hyperparameters, and hardware configuration.
@@ A.1 Proofs of Theorems
The derivative of the log-likelihood function in eq. 4 can be expressed as
@@
Following the same steps of the proof for theorem 5.2, we can show that $\mathbb{E}_{X}\left[\sigma_{\beta^{*}}(X)\right]=\mathbb{E}_{Y}\left[Y\right]$ and ${\mathbb{E}_{X}\left[X_{\left[i\right]}\;\sigma_{\beta^{*}}(X)\right]=\mathbb{E}_{Y}\left[Y\;\mathbb{E}_{X\mid Y}\left[X_{\left[i\right]}\right]\right]}$.
@@ A.2 Connections to Maximum Mean Discrepancy
[…]
@@
The formulation presented in section 5 shares some similarities with Maximum Mean Discrepancy (MMD) used by VFAE. Indeed, MMD states that given two distributions, they are equal as long as the maximum distance between their expected value under a transformation $f$ is zero.
Thus, if $\operatorname{MMD}_{\mathcal{H}}\left(\mathbb{P}_{X},\mathbb{P}_{Y}\right)=0$ then $\mathbb{P}_{X}=\mathbb{P}_{Y}$. This is indeed exploited by Adversarial Learning, which tries to find $f$ by means of a min-max optimization. However, as shown by [2], Adversarial Learning fails as finding such $f$ is extremely hard, due to the fact that the two distributions are constantly changing based on the current $f$. Therefore, if $\operatorname{MMD}_{\mathcal{H}}\left(\mathbb{P}_{X},\mathbb{P}_{Y}\right)\neq 0$, no provable guarantee on the fairness can be given. Instead, if $f$ is in the family of LR, theorem 5.3 shows that the second moment bounds the predictive power of such a representation.
@@ A.3 Matching All Moments
The approaches presented in sections 4 and 5 can be reconducted to matching the moment generating function, or directly the moments of a distribution. Indeed, matching all moments of two distributions with bounded support ensures that they are equal almost everywhere. Even though it is trivial to force the support of $Z$ to be bounded, directly applying moment matching to the distributions $\mathbb{P}_{Z\mid S}$ might be impractical, as all moments are needed to characterize a distribution, which is computationally unfeasible. Furthermore, truncating to the $k$-th moment leads to a biased estimation, and, in the context of this work, potential leakage of the sensitive attribute $S$.
@@
Instead, the proposed approaches presented in sections 4 and 5 use a more principled and computationally stable way of matching two distributions.
@@ A.5 Additional empirical evaluations and ablations
In table 3(a) we compare our proposed approaches to other baselines on the Adult dataset, with the same setup as in [37]. Given that no distinction is made on the family of functions used for the evaluation, we report the Adv. Accuracy of FmCF and FmSS approaches using a deep and wide MLP. The same holds true for the results reported in table 3(b), where we compare out methods to the same setup from [47]
@@
In figs. 8, 6, 7 and 5, we report the tradeoffs between adversarial balanced accuracy and task accuracy. We evaluate FmCF section 4 against an adversarial MLP as reported in section A.7. Instead, we evaluate FmSS section 5 both against an MLP and against a LR. In particular, for all images, in black is reported the performance of Fair Normalizing Flows [2], while in blue is reported the performance using 1 dimension, in orange 2 dimensions, and in green 3 dimensions. We observe that the lower the dimension, the easier it is for the penalties to be effective, giving surprising performances for extremely low-dimensional latent spaces. Notably, such a characteristic is not limiting for the downstream task, as even in the lowest-dimensional setting, both approaches outperform all the baselines.
@@
Given that all benchmarks considered thus far pertain to classification tasks, we now present a simple example demonstrating the applicability of FmCF to alternative problem settings. Specifically, in fig. 9, a conditional convolutional autoencoder is employed to learn a latent representation of the MNIST handwritten digits dataset. Subsequently, the loss function described in algorithm 1 is utilized to penalize the divergence between $\mathbb{P}_{Z\mid S}$ and a Gaussian distribution. When this condition is satisfied, it ensures that $Z\perp\!\!\!\perp S$, indicating that the latent representation $Z$ encodes information about $X$ that is independent of $S$. Consequently, when the model is prompted to reconstruct an input with a different label, it retains the essential characteristics of the original input while generating a digit consistent with the newly specified label.
@@
For more details on the datasets, please refer to section A.6.

### v2

@@ Appendix A Appendix
  • Section˜A.1: Complete derivations for Theorems 5.1, 5.2, and 5.3.
  • Section˜A.2: Analysis of MMD, its relation to characteristic‐function distances, and links to adversarial representation learning.
  • Section˜A.3: Discussion of full moment‐matching, the Russian‐Roulette estimator for unbiased series truncation, and practical considerations.
  • Section˜A.4: Formal definitions and commentary on Demographic Parity, Equalized Odds, and other group‐fairness criteria.
  • Section˜A.5: Extended comparisons against baselines, Pareto‐front ablations over latent dimensions, and robustness checks.
  • Section˜A.6: Detailed descriptions of all benchmark datasets, preprocessing pipelines, and protected attribute encoding.
  • Section˜A.7: Implementation specifics, optimizer settings, hyperparameters, and hardware configuration.
@@ A.1 Proofs of Theorems
The derivative of the log-likelihood function in eq.˜4 can be expressed as
@@
Following the same steps of the proof for theorem˜5.2, we can show that $\mathbb{E}_{X}\left[\sigma_{\beta^{*}}(X)\right]=\mathbb{E}_{Y}\left[Y\right]$ and $\mathbb{E}_{X}\left[X_{\left[i\right]}\;\sigma_{\beta^{*}}(X)\right]=\mathbb{E}_{Y}\left[Y\;\mathbb{E}_{X\mid Y}\left[X_{\left[i\right]}\right]\right]$}.
@@ A.2 Connections to Maximum Mean Discrepancy
$\operatorname{MMD}_{\mathcal{H}}^{2}\left(\mathbb{P}_{X},\mathbb{P}_{Y}\right)=\int_{\mathbb{R}^{d}}\left\lvert\varphi_{X}(t)-\varphi_{Y}(t)\right\rvert^{2}w(t)dt.$ (37)
@@
The formulation presented in section˜5 shares some similarities with Maximum Mean Discrepancy (MMD) used by VFAE. Indeed, MMD states that given two distributions, they are equal as long as the maximum distance between their expected value under a transformation $f$ is zero.
Thus, if $\operatorname{MMD}_{\mathcal{H}}\left(\mathbb{P}_{X},\mathbb{P}_{Y}\right)=0$ then $\mathbb{P}_{X}=\mathbb{P}_{Y}$. This is indeed exploited by Adversarial Learning, which tries to find $f$ by means of a min-max optimization. However, as shown by [2], Adversarial Learning fails as finding such $f$ is extremely hard, due to the fact that the two distributions are constantly changing based on the current $f$. Therefore, if $\operatorname{MMD}_{\mathcal{H}}\left(\mathbb{P}_{X},\mathbb{P}_{Y}\right)\neq 0$, no provable guarantee on the fairness can be given. Instead, if $f$ is in the family of LR, theorem˜5.3 shows that the second moment bounds the predictive power of such a representation.
@@ A.3 Matching All Moments
The approaches presented in sections˜4 and 5 can be reconducted to matching the moment generating function, or directly the moments of a distribution. Indeed, matching all moments of two distributions with bounded support ensures that they are equal almost everywhere. Even though it is trivial to force the support of $Z$ to be bounded, directly applying moment matching to the distributions $\mathbb{P}_{Z\mid S}$ might be impractical, as all moments are needed to characterize a distribution, which is computationally unfeasible. Furthermore, truncating to the $k$-th moment leads to a biased estimation, and, in the context of this work, potential leakage of the sensitive attribute $S$.
@@
Instead, the proposed approaches presented in sections˜4 and 5 use a more principled and computationally stable way of matching two distributions.
@@ A.5 Additional empirical evaluations and ablations
In table˜3(a) we compare our proposed approaches to other baselines on the Adult dataset, with the same setup as in [37]. Given that no distinction is made on the family of functions used for the evaluation, we report the Adv. Accuracy of FmCF and FmSS approaches using a deep and wide MLP. The same holds true for the results reported in table˜3(b), where we compare out methods to the same setup from [47]
@@
In figs.˜8, 6, 7 and 5, we report the tradeoffs between adversarial balanced accuracy and task accuracy. We evaluate FmCF section˜4 against an adversarial MLP as reported in section˜A.7. Instead, we evaluate FmSS section˜5 both against an MLP and against a LR. In particular, for all images, in black is reported the performance of Fair Normalizing Flows [2], while in blue is reported the performance using 1 dimension, in orange 2 dimensions, and in green 3 dimensions. We observe that the lower the dimension, the easier it is for the penalties to be effective, giving surprising performances for extremely low-dimensional latent spaces. Notably, such a characteristic is not limiting for the downstream task, as even in the lowest-dimensional setting, both approaches outperform all the baselines.
@@
Furthermore, we evaluated the effects of batch size on the reliability of the CFD estimation. As confirmed by table˜4, larger batches yield more stable and reliable CFD estimates. The variance of the estimates decreases consistently with increasing batch size, aligning with the expected theoretical scaling of $O(\frac{1}{n})$.
Table 4: Standard deviation of FmCF Penalty estimator.
Batch Size Adult German Compas Health
8 0.1150 0.1252 0.0924 0.1004
16 0.0596 0.0854 0.0361 0.0422
32 0.0164 0.0172 0.0129 0.0121
64 0.0063 0.0056 0.0052 0.0054
128 0.0026 0.0024 0.0022 0.0027
256 0.0016 0.0004 0.0010 0.0015
512 0.0007 0.0000 0.0007 0.0009
Given that all benchmarks considered thus far pertain to classification tasks, we now present a simple example demonstrating the applicability of FmCF to alternative problem settings. Specifically, in fig.˜9, a conditional convolutional autoencoder is employed to learn a latent representation of the MNIST handwritten digits dataset. Subsequently, the loss function described in algorithm˜1 is utilized to penalize the divergence between $\mathbb{P}_{Z\mid S}$ and a Gaussian distribution. When this condition is satisfied, it ensures that $Z\perp\!\!\!\perp S$, indicating that the latent representation $Z$ encodes information about $X$ that is independent of $S$. Consequently, when the model is prompted to reconstruct an input with a different label, it retains the essential characteristics of the original input while generating a digit consistent with the newly specified label.
@@
For more details on the datasets, please refer to section˜A.6.